P3-29. Contribution Income Statement, Cost-Volume-Profit Graph, and Taxes New York Tours (NYT) provides daily sightseeing tours that include transportation, admission to se- lected attractions, and lunch. Ticket prices are S85 each. During June 2017, NYT provided 3,000 tours. Variable Costs per Customer Fixed Costs per Month $25,000 15,000 Lunch Overhead Selling and administrative Total Selling and administration 17 10 8 $65 Total NYT is subject to an income tax rate of 40 percent.

Required

a. Prepare a contribution income statement for June.

b. Determine NYT’s monthly break-even point in units.

c. Determine NYT’s margin of safety for June 2017.

d. Determine the unit sales required for a monthly after-tax profit of $15,000.

e. Prepare a cost-volume-profit graph. Label the horizontal axis in units with a maximum value of

4,000. Label the vertical in dollars with a maximum value of $300,000. Draw a vertical line on

the graph for the current (3,000) unit level and label total variable costs, total fixed costs, and total

before-tax profits at 3,000 units.
Answer #1

Answer

  • All working forms part of the answer
  • As per guidelines, we have to answer first four parts only when a question has subparts. However, I tried the last part also.
  • Answer (a)

Service Revenues

[3000 x $85]

$255,000

(-) Variable cost

[3000 x $65]

$195,000

Contribution margin

[3000 x $(85-65)]

$60,000

(-) Total Fixed cost

$40,000

Net Income

$20,000

  • Answer (b)

A

Total Contribution margin

$60000

B

Total Units

3000

C=A/B

Unit Contribution margin

$20

D

Total Fixed cost

$40000

E=D/C

Break Even point in Units

2000

  • Answer (c)

A

Break Even point in Units

2000

B

Total tour

3000

C=B-A

Margin of Safety in Units

1000

D

Price per unit

$85

E=C x D

Margin of Safety in $

$85,000

  • Answer (d)

A

Target profits

$15000

B

Total Fixed cost

$40000

C=A+B

Total contribution margin required for target profits

$55,000

D

Unit Contribution margin

$20

E=C/D

Units sale required for target profit

2750

  • Answer (e)

300000 260000 oSL 250000 200000 195000 178750 Total variable cost Total Fixed Cost , Total Profits 130000 100000 50000 40000